# Probabilistic Consistency in Machine Learning and Its Connection to Uncertainty Quantification

### v1

@@ 4.2.3 Binary Classifiers Revisited
Let $\hat{C}:\Gamma\times{{\rm X}^{2}}\to\{1,2\}$ be a self-consistent binary classifier, where $\Gamma\subset\mathbb{R}^{n}$ is a compact set for $n\geq 1$. If $R_{i,j}$ is positive, bounded, and has the properties that $0<R_{i,j}<1$ on a set of finite Lebesgue measure, and moreover, $1<R_{i,j}<M<\infty$ on a set of finite Lebesgue measure, then there exist density functions $P_{1}(r)$ and $P_{2}(r)$ for which $\hat{C}$ is the Bayes classifier.
@@
We proceed by construction. Let $\Lambda_{0}=\{r:R_{1,2}=1\}$, $\Lambda_{1}=\{r:R_{1,2}(r)<1\}$, and $\Lambda_{2}=\{r:R_{1,2}(r)>1\}$. Let $\hat{P}_{1}(r)=\alpha_{m}$ for $r\in\Lambda_{m}$, where $\alpha_{m}$ are constants and $m=0,1,2$. Normalization of $\hat{P}_{2}(r)$ implies the identity
$\alpha_{1}\int_{\Lambda_{1}}\!\!{\rm d}r\,\,(R_{1,2}(r)-1)+\alpha_{2}\int_{\Lambda_{2}}\!\!{\rm d}r\,\,(R_{1,2}(r)-1)=0.$ (35)
By compactness, both integrals are finite, so that Eq. (35) is valid linear equation, which also implies that $\alpha_{1}$ and $\alpha_{2}$ have the same sign. Moreover, normalization of $\hat{P}_{1}$ implies
$\sum_{m=0}^{2}\mu(\Lambda_{m})=1$ (36)
where $\mu(\Lambda_{m})$ is the Lebesgue measure of the $\Lambda_{m}$. This leads to a system of equations for $\alpha_{m}$, which yields a normalized and strictly positive $\hat{P}_{1}(r)$. Moreover, $\hat{P}_{2}(r)=R_{1,2}(r)\hat{P}_{1}(r)$ is normalized. Thus, there exist density functions for which $\hat{C}:\Gamma\times{{\rm X}^{2}}\to\{1,2\}$ is the Bayes classifier for all test prevalence values $\chi\in{{\rm X}^{2}}$.
@@ 5.0.1 Examples of Induced Density Ratios
[…]
@@
where ??? stands for an indeterminate form whose value we cannot anticipate.
@@
where the elements of the square matrix ${\boldsymbol{\rm B}}$ satisfy Eq. (26), $\boldsymbol{\infty}$ is a matrix of whose elements are all $\infty$, and ${\boldsymbol{0}}$ is a matrix whose elements are all zero.
@@ 5.1 Illustration of Construction 1: Gaussian Distributions
where $\Sigma_{j}$ is a $\times 2$ 2 covariance matrix, and $\mu_{j,x}$ and $\mu_{j,y}$ are fixed parameters. For $j\in\{1,2,3\}$, these are chosen to be
@@ 5.2.1 MNIST Dataset and Neural Network Architecture
The full dataset is comprised of 60,000 training images of handwritten numerals ‘0’ to ‘9’, along with 10,000 validation images. (We use single-quotes to distinguish handwritten numerals from numbers.) Here we consider the restricted three-class problem of identifying the numerals ‘1’, ‘2’, and ‘7’, which can be difficult to distinguish visually. We let $\omega$ denote a handwritten numeral, $r(\omega)$ the corresponding grid of $\times 28$ 28 pixel values associated with its image, and $C(\omega)=k\in\{1,2,3\}$ the true class of the image under the mapping 11$\mapsto$1 ↦ ‘1’, 22$\mapsto$2 ↦ ‘2’, and 33$\mapsto$3 ↦ ‘7’. Properties of the training and test data are listed in Table 1. Note that $s_{k}$ denotes the number of samples in the $k$th class.
@@
Figure 4: Schematic of the neural network architectures used in this work. The input is a $\times 28$ 28 pixel image with 256 grayscale levels. This feeds into a convolutional layer with a convolution width of 555 pixels and zero-padding, which ensures that the output (large yellow rectangle with smaller orange ractangle) is the same size as the input. Next we apply a normalization layer (blue), followed by a rectified linear unit (ReLu) layer (light orange). The output of this is fed to another set of layers with an identical structure, followed by another. The end-result is fed into a softmax layer. Reference [70] was used to generate this figure.
@@
A key problem with Eq. (44) is that it is not differentiable. As such, backpropagation is not possible [41], so that optimization methods such as stochastic gradient descent cannot be used [41]. One solution to this is to consider a homotopy-type approach wherein we consider a sequence of objective functions that converge Eq. (44) as a scaling parameter $\sigma\to 0$. See Refs. [72, 73, 74, 75, 76] for general background on homotopy approaches, as well as [42, 71, 77] for uses and variations of this method in the context of ML. Here we consider the homotopy objective function as a composition of several functions:
@@ 5.2.2 Self-Consistency
Figure 5: Illustration of feasibility and empirical self-consistency in a three-class setting. The horizontal and vertical axes correspond to values of $R_{1,2}$ and $R_{2,7}$, while the hyperbolas are curves of constant $R_{1,2}R_{2,7}$. The shaded box shows an example of binning uncertainty in the values associated with $R_{1,2}$ and $R_{1,7}$. For a fixed value of $r$, the triple $\Psi(r)=(R_{1,2},R_{2,7},R_{1,7})$ is only consistent if $R_{2,7}=R_{1,2}R_{2,7}$. If a monotone classifier indicates that $R_{1,7}$ is bounded by some values $[R_{1,7}^{\ell},R_{1,7}^{h}]$, then it is consistent if and only if the domain bounded by the corresponding hyperbolas intersects the box. Moreover, that intersection is the feasible set. In the examples that follow, we define $u(r)$ to be the arithmetic mean of the inherent uncertainties computed on the vertices of such domains.
@@ 5.2.3 Numerical Results
Table 2 summarizes the outcomes of our numerical experiments on the MNIST data. After training the family of pairwise classifiers and computing bounds on the density ratios, our analysis rejected 95 training points as inconsistent with the law of total probability. Figure 6 shows all 95 of these images. Such data are inputs for which we cannot meaningfully assign a class because the algorithm yields results that contradict the law of total probability; recall Eq. (25). It is comforting that many of these images bear no resemblance whatsoever to any of the numerals ‘1’, ‘2’, or ‘7’. This suggests that the concept of consistency is in fact useful for identifying data that cannot be analyzed by the NN. Figures 7 and 8 show sample images from the training set that have low and high probabilities of being classified correctly.
@@ 6.2 Considerations for Optimization: Homotopy Methods, H-Consistency, and Cross Entropy
×$\mathcal{E}_{\times}(\Pi_{tr},\phi)$ $=\sum_{k=1}^{K}\frac{s_{k}}{s_{k}N}\sum_{j=1}^{s_{k}}-\log(y_{j,k,k}(\phi))=\sum_{k=1}^{K}\frac{\bar{q}_{k}}{s_{k}}\sum_{j=1}^{s_{k}}-\log(y_{j,k,k}(\phi)),$ (50)

### v2

@@ 4.2.3 Binary Classifiers Revisited
Let $\hat{C}:\Gamma\times{{\rm X}^{2}}\to\{1,2\}$ be a self-consistent binary classifier, where $\Gamma\subset\mathbb{R}^{n}$ is a compact set for $n\geq 1$. If the induced density ratio $\hat{R}_{1,2}$ is positive, bounded, and has the properties that $0<\hat{R}_{1,2}<1$ on a set of finite Lebesgue measure, and moreover, $1<\hat{R}_{1,2}<M<\infty$ on a set of finite Lebesgue measure, then there exist density functions $P_{1}(r)$ and $P_{2}(r)$ for which $\hat{C}$ is the Bayes classifier.
@@
We proceed by construction. Let $\Lambda_{0}=\{r:\hat{R}_{1,2}(r)=1\}$, $\Lambda_{1}=\{r:\hat{R}_{1,2}(r)<1\}$, and $\Lambda_{2}=\{r:\hat{R}_{1,2}(r)>1\}$. Let $\hat{P}_{1}(r)=\alpha_{m}$ for $r\in\Lambda_{m}$, where $\alpha_{m}$ are constants and $m=0,1,2$. Normalization of $\hat{P}_{2}(r)$ implies the identity
$\alpha_{1}\int_{\Lambda_{1}}\!\!{\rm d}r\,\,(\hat{R}_{1,2}(r)-1)+\alpha_{2}\int_{\Lambda_{2}}\!\!{\rm d}r\,\,(\hat{R}_{1,2}(r)-1)=0.$ (35)
By compactness, both integrals are finite, so that Eq. (35) is valid linear equation, and it also implies that $\alpha_{1}$ and $\alpha_{2}$ have the same sign. Moreover, normalization of $\hat{P}_{1}$ implies
$\sum_{m=0}^{2}\alpha_{m}\mu(\Lambda_{m})=1$ (36)
where $\mu(\Lambda_{m})$ is the Lebesgue measure of the $\Lambda_{m}$. This leads to a system of equations for $\alpha_{m}$, which yields a normalized and strictly positive $\hat{P}_{1}(r)$. Moreover, $\hat{P}_{2}(r)=R_{1,2}(r)\hat{P}_{1}(r)$ is normalized as a result of Eq. (35). Thus, there exist density functions for which $\hat{C}:\Gamma\times{{\rm X}^{2}}\to\{1,2\}$ is the Bayes classifier for all test prevalence values $\chi\in{{\rm X}^{2}}$.
@@ 5.0.1 Examples of Induced Density Ratios
We begin with a series of three- and four-class examples that illustrate certain properties of Lemma 2 and Proposition 3. The goal of these examples is to address the practical issue of identifying the sets $W$ and $V$, especially in a multiclass setting with large $K$. Here we assume that the $\hat{R}_{i,j}$ are both known exactly and self-consistent, focusing on the fact that these must be evaluated pointwise for each $r$ in an empirical test population. In this context, it is convenient to define an induced density matrix $\hat{\boldsymbol{\rm R}}$ whose $(i,j)$th element is $\hat{R}_{i,j}$. We temporarily omit any dependence on the variable $r$, since this can be treated as fixed.
@@
where ?? stands for an indeterminate form whose value we cannot anticipate.
@@
[…]
@@ 5.1 Illustration of Construction 1: Gaussian Distributions
where $\Sigma_{j}$ is a $2\times 2$ covariance matrix, and $\mu_{j,x}$ and $\mu_{j,y}$ are fixed parameters. For $j\in\{1,2,3\}$, these are chosen to be
@@ 5.2.1 MNIST Dataset and Neural Network Architecture
The full dataset is comprised of 60,000 training images of handwritten numerals ‘0’ to ‘9’, along with 10,000 validation images. (We use single-quotes to distinguish handwritten numerals from numbers.) Here we consider the restricted three-class problem of identifying the numerals ‘1’, ‘2’, and ‘7’, which can be difficult to distinguish visually. We let $\omega$ denote a handwritten numeral, $r(\omega)$ the corresponding grid of $28\times 28$ pixel values associated with its image, and $C(\omega)=k\in\{1,2,3\}$ the true class of the image under the mapping $1\mapsto$ ‘1’, $2\mapsto$ ‘2’, and $3\mapsto$ ‘7’. Properties of the training and test data are listed in Table 1. Note that $s_{k}$ denotes the number of samples in the $k$th class.
@@
Figure 4: Schematic of the neural network architectures used in this work. The input is a $28\times 28$ pixel image with 256 grayscale levels. This feeds into a convolutional layer with a convolution width of 55 pixels and zero-padding, which ensures that the output (large yellow rectangle with smaller orange ractangle) is the same size as the input. Next we apply a normalization layer (blue), followed by a rectified linear unit (ReLu) layer (light orange). The output of this is fed to another set of layers with an identical structure, followed by another. The end-result is fed into a softmax layer. Reference [70] was used to generate this figure.
@@
A key problem with Eq. (44) is that it is not differentiable. As a result, backpropagation is not possible [41], so that optimization methods such as stochastic gradient descent cannot be used [41]. One solution to this is to consider a homotopy-type approach wherein we consider a sequence of objective functions that converge Eq. (44) as a scaling parameter $\sigma\to 0$. See Refs. [72, 73, 74, 75, 76] for general background on homotopy approaches, as well as [42, 71, 77] for uses and variations of this method in the context of ML. Here we consider the homotopy objective function as a composition of several functions:
@@ 5.2.2 Self-Consistency
Figure 5: Illustration of feasibility and empirical self-consistency in a three-class setting. The horizontal and vertical axes correspond to values of $R_{1,2}$ and $R_{2,7}$, while the hyperbolas are curves of constant $R_{1,2}R_{2,7}$. The shaded box shows an example of binning uncertainty in the values associated with $R_{1,2}$ and $R_{1,7}$. For a fixed value of $r$, the triple $\Psi(r)=(R_{1,2},R_{2,7},R_{1,7})$ is only consistent if $R_{2,7}=R_{1,2}R_{2,7}$. If a monotone classifier indicates that $R_{1,7}$ is bounded by some values $R_{1,7}^{\ell},R_{1,7}^{h}]$, then it is consistent if and only if the domain bounded by the corresponding hyperbolas intersects the box. Moreover, that intersection is the feasible set. In the examples that follow, we define $u(r)$ to be the arithmetic mean of the inherent uncertainties computed on the vertices of such domains.
@@ 5.2.3 Numerical Results
Table 2 summarizes the outcomes of our numerical experiments on the MNIST data. After training the family of pairwise classifiers and computing bounds on the density ratios, our analysis rejected 95 training points as inconsistent with the law of total probability. Figure 6 shows all 95 of these images. Such data are inputs for which we cannot meaningfully assign a class because the algorithm yields results that contradict the law of total probability; recall Eq. (25). It is comforting that many of these images bear no resemblance whatsoever to any of the numerals ‘1’, ‘2’, or ‘7’. This suggests that the concept of self-consistency is in fact useful for identifying data that cannot be analyzed by the NN. Figures 7 and 8 show sample images from the training set that have low and high probabilities of being classified correctly.
@@ 6.2 Considerations for Optimization: Homotopy Methods, H-Consistency, and Cross Entropy
$\mathcal{E}_{\times}(\Pi_{tr},\phi)$ $=\sum_{k=1}^{K}\frac{s_{k}}{s_{k}N}\sum_{j=1}^{s_{k}}-\log(y_{j,k,k}(\phi))=\sum_{k=1}^{K}\frac{\bar{q}_{k}}{s_{k}}\sum_{j=1}^{s_{k}}-\log(y_{j,k,k}(\phi)),$ (50)